OpenStudy (anonymous):

Given: angle 1 = angle 2 If AC = 3, BC = 5, and AB = 7, find AD: A. 2 5/8 B. 4 1/5 C. 4 3/8 D. 11 2/3

OpenStudy (anonymous):

what is 1 and 2?

OpenStudy (anonymous):

angle 1= angle 2.

OpenStudy (anonymous):

hotty

OpenStudy (anonymous):

can you see two similar triangles?

OpenStudy (anonymous):

AC/AD = BD/BC AD+BD=AB

OpenStudy (anonymous):

@Diana691 can you solve now?

OpenStudy (anonymous):

okay but AD and BD are unknown.

OpenStudy (anonymous):

Yes we have two variables and two equation

OpenStudy (anonymous):

AD+DB= AB

OpenStudy (anonymous):

can you show me step by step. please

OpenStudy (anonymous):

assume AD =x then DB = 7-x solve the quadratic equation

OpenStudy (anonymous):

let us do it together here. if you are stuck I will help you!

OpenStudy (anonymous):

@Diana691 put the values in second equation and form the equation

OpenStudy (anonymous):

okay -x squared =7x?

OpenStudy (anonymous):

3*5 = x*(7-x)

OpenStudy (anonymous):

3x=35-5

OpenStudy (anonymous):

NOoooo

OpenStudy (anonymous):

what am i doing wrong? ;/

OpenStudy (anonymous):

can you make equation from this 3*5 = x*(7-x) * means multiply.

OpenStudy (anonymous):

x squared -7x+15

OpenStudy (anonymous):

can you help me @GuttierreOliveira

OpenStudy (anonymous):

@nikato

OpenStudy (anonymous):

Try finding the angles.

OpenStudy (nikato):

If u still need help, can u tell me if there is any more given info like if any r right angles

OpenStudy (anonymous):

Hi @Diana691 , do not know if you've managed to make the exercise more here's my explanation lol this is basically the same thing as the others ! when he says that the angle 1 = angle 2 means that the legs assumed the same values ​​and the same proportion (not sure if this definition is correct , but I think it is ) . eg if the triangle which has 1 angle is 30 ° and a hypotenuse with value 3, where the line AC , you can find the AD line using the law of sin ! Now if you performed the same operation on the triangle that takes the angle 2 has already been defined (if angle 1 = angle 2 then angle 2 = 30 ° ) using the same values ​​above , you agree that the result would be the same ? However , if we change the value of a leg will change the result , but in the same proportion :) and change the angle values ​​also changed , (because sin 30 ° = 0.50 and sin 60 ° = 0.8660 ... . ) As we do not have the value of the angle, we must do what we had already done ! AC = 3 BC = 5 AB = 7 - AD AD = x AC/BC = AD/AB 3/5 = x/7 - x  ( AD = X then I replace here) 5x = 21 - 3x 8x = 21 x = 21 / 8 or 2,62... I do not quite understand the answers but I think "A" is because 5/8 = 0.625 be added + 2 = 2,625 more if it is 25/8 . I do not know how to do ! :)